OpenStudy (anonymous):

Michele Gabrielle borrowed $6,000 for 3 months and paid $82.50 in simple interest on the loan. Find the annual simple interest rate that Michele paid on the loan. (Round your answer to the nearest tenth of a percent.)

OpenStudy (espex):

For simple interest use: I=p*r*t

OpenStudy (anonymous):

i just have trouble rounding the number?

OpenStudy (espex):

i = 82.50, p = 6000, t = 3/12 \[\frac{82.50}{6000 * .25} = rate\] Okay, what was your answer before rounding?

OpenStudy (anonymous):

0.00055

OpenStudy (anonymous):

soeey it 0.055

OpenStudy (anonymous):

sorry

OpenStudy (espex):

Unless I did my math wrong you are a couple extra decimals off. For mine I got 0.05.5

OpenStudy (espex):

Okay, so 0.055, times 100 since they ask for percent, give you 5.5% Now just round to the nearest 10th.

OpenStudy (anonymous):

6.0?

OpenStudy (espex):

That is how I would interpret this problem. :)

OpenStudy (anonymous):

it didn't take that answer :(

OpenStudy (espex):

Personally I would have left it at 5.5%, but since it asks you to round then 6.0 seems to be the correct answer.

OpenStudy (anonymous):

well it took 5.5 Thanks Much :)

OpenStudy (espex):

You are welcome. :)